OpenStudy (anonymous):

Assuming that the pentagon is regular, what is the area of the shaded region below? Use your calculator and round your final answer to the nearest hundredth.

OpenStudy (anonymous):

shaded region?

OpenStudy (anonymous):

Area of pentagon - area of square

OpenStudy (anonymous):

and then what do i do?

OpenStudy (anonymous):

regular pentagon means all sides are equal \[area = \frac{ 1 }{ 4 }\sqrt{5(5+2\sqrt{5})}* side^2\]

OpenStudy (anonymous):

side = 9

OpenStudy (anonymous):

so the answer is 58.36 ?

OpenStudy (anonymous):

Correct. :)